OpenStudy (anonymous):

You have challenged four other students to come up with expressions equal to x^2. Their submissions are below.

OpenStudy (anonymous):

Explain the steps in simplifying the expression of each student and identify who is correct.

OpenStudy (anonymous):

@BeckieBird @shamil98 @pitamar @AllTehMaffs Thanks guys :D

OpenStudy (anonymous):

o.o uhm ok

OpenStudy (anonymous):

I would say Kelly \[ \sqrt[7]{ x^3 \cdot x^5 \cdot x^6 } = \sqrt[7]{ x^{3+5+6} } = \sqrt[7]{x^{14}} =\\ = \sqrt[7]{x^7 \cdot x^7} = x \cdot x = x^2 \]

OpenStudy (anonymous):

Thanks! Can you elaborate more though, can you explain it?

OpenStudy (anonymous):

agreed; Kelly You can add the exponents under the 7th root normally, and then divide by 7 \[\sqrt[7]{x ^{3}x^{5}x^{6}}=x ^{\frac{ (3+5+6) }{ 7 }} = x^{2}\] Mike's is also right : \[\frac{ x ^{\frac{ 8 }{ 5 }}x ^{\frac{ 4 }{ 5 }} }{ x ^{\frac{ 2 }{ 5 }} } = x ^{\frac{ 8+4-2 }{ 5 }}=x^{2}\] ....

OpenStudy (anonymous):

Hmm, I didn't even look at Mike, good job hehe.. For some reason I thought there is one solution

OpenStudy (anonymous):

Kim's works too; (1+2+3+4)/5 =2 Scott seems wonky though. His just gives the square root of x

OpenStudy (anonymous):

That crazy Scott

OpenStudy (anonymous):

:) thanks